Assuming an interest rate of 6.6 percent, what is the value of the following cash flows five years from today? Year Cash Flow 1 $ 3,615 2 4,715 3 5,705 4 7,000
Answer #1

Future value = CF(1+r)n

Future value = CF (1+r)4 +CF (1+r)3 +CF (1+r)2 +CF (1+r)1

Future value= 3615 (1+.066)4 + 4715(1.066)3 + 5705 (1+.066)2 + 7000(1+.066)= 4668.07+ 5711.54+ 6482.91+ 7462=24324.52

CF means cash flow,

r means rate of interest

n mean number of period